OpenStudy (anonymous):

Simplify the expression. √6(√33+7) A. √39+7√6 B. √198+7 C. 3√22+7√6 D. √198+7√6

OpenStudy (anonymous):

@experimentX please help

OpenStudy (anonymous):

I got C

OpenStudy (jhannybean):

oh my goodness.bad mistake.

OpenStudy (jhannybean):

it's not sqrt{37} it's sqrt{33} -_-

OpenStudy (anonymous):

so the answer is not c then

OpenStudy (jhannybean):

\[\large \sqrt{6}(\sqrt{33}+7)\]\[\large \sqrt{a}*\sqrt{b} \implies \sqrt{6}*\sqrt{33}=\sqrt{198}\]\[\large \sqrt{6}*7 \implies 7\sqrt{6}\]altogether \[\large \sqrt{6}(\sqrt{33}+7)= \sqrt{198}+7\sqrt{6}\]

OpenStudy (anonymous):

o ok thanks so much

OpenStudy (jhannybean):

Sorry about that... i wrote \(\sqrt{37} \) instead of \(\sqrt{33}\)

OpenStudy (anonymous):

so you saying it was D and that is not write

OpenStudy (anonymous):

right sorry

OpenStudy (anonymous):

it came up wrong

OpenStudy (jhannybean):

it came up wrong? how come?

OpenStudy (saifoo.khan):

OMG! Jaan baji made a mistake!! =D

OpenStudy (anonymous):

\[\Huge \color{Red}{\bf \bigstar}\]

OpenStudy (jhannybean):

You are right @Tabbiejack .\[\large \sqrt{198} = \sqrt{18*11} = \sqrt{9*2*11} = \sqrt{3*3*2*11} = 3\sqrt{22}\] So this would be \[\large 3 \sqrt{22}+7\sqrt{6}\] goodjob.